3x/4+5=2x/3: 3x/4+5=2x/3
Answer by checkley77(3639)

3x/4+5=2x/3
3X/4-2X/3=-5
(3*3X-4*2X)/4*3=-5
(9X-8X)/12=-5
X/12=-5
X=-5*12
X=-60
PROOF:
3*-60/4+5=2*-60/3
-180/4+5=-120/3
(-180+4*5)/4=-40
(-180+20)/4=-40
-160/4=-40
-40=-40